Point B has coordinates (-11,11)and lies on line BS, which has a slope of -5/6.
Find the equation of line BS in slope-intercept form.

Answer by richwmiller(17219)

y=mx+b
we know x=-11 we know y =11 we know m=-5/6
11=-5/6*(-11)+b
11=55/6+b
66/6-55/6=b
11/6=b
plug them in and find b and plug into
y=-5/6x+b
y=-5/6x+11/6